OpenStudy (anonymous):

Evaluate the following expression using the values given. Find 2x2 - 2z4 + y2 - x2 + z4 if x = -4, y = 3, and z = 2.

OpenStudy (sleepyjess):

\(2x^2-2z^4+y^2-x^2+z^4=2(-4)^2-2(2)^4+3^2-(-4)^2+2^4\) Combine like terms \(-4^2-2^4+3^2\) \(-4^2=16\) \(-2^4=16\) \(3^2=9\) \(16+16+9\) now solve